### v2

@@ 4.4 Ablative Study and Analysis
Ablation of loss components in FDM. To verify the effectiveness of the specific loss components within the FDM, we conduct ablation studies on the identity loss $\mathcal{L}_{I}$ and reconstruction loss $\mathcal{L}_{r}$. These experiments are performed based on the best-performing Qwen2.5-VL-7B model, and evaluated on the two challenging WDF and DFDC datasets under the fuse-dataset setup. The results are shown in Table 5. It could be observed that the performance improves on the WDF dataset after introducing the $\mathcal{L}_{I}$. This indicates that explicit identity supervision helps MARE to decouple identity information from forgery traces features, reducing the interference of identity signals on authenticity detection. Besides, MARE achieves the best performance on both WDF and DFDC after the introduction of $\mathcal{L}_{r}$, which demonstrates that $\mathcal{L}_{r}$ imposes a crucial information constraint that ensures the stability of the disentanglement architecture. Therefore, both $\mathcal{L}_{I}$ and $\mathcal{L}_{r}$ help to improve the ability of MARE in capturing intrinsic forgery traces features.
@@ 5 Conclusion
Appendix A More experimental analysis

### v3

@@ 4.4 Ablative Study and Analysis
Ablation of loss components in FDM. To verify the effectiveness of the specific loss components within the FDM, we conduct ablation studies on the identity loss $\mathcal{L}_{I}$ and reconstruction loss $\mathcal{L}_{r}$. These experiments are performed based on the best-performing Qwen2.5-VL-7B model, and evaluated on the two challenging WDF and DFDC datasets under the fuse-dataset setup. The results are shown in Table 5. It could be observed that the performance improves on the WDF dataset after introducing the $\mathcal{L}_{I}$. This indicates that explicit identity supervision helps MARE to decouple identity information from forgery traces features, reducing the interference of identity signals on authenticity identification. Besides, MARE achieves the best performance on both WDF and DFDC after the introduction of $\mathcal{L}_{r}$, which demonstrates that $\mathcal{L}_{r}$ imposes a crucial information constraint that ensures the stability of the disentanglement architecture. Therefore, both $\mathcal{L}_{I}$ and $\mathcal{L}_{r}$ help to improve the ability of MARE in capturing intrinsic forgery traces features.
@@ 5 Conclusion
Appendix A Additional Experimental Results
@@ Appendix B Facial Region Set
Appendix C Limitation and Future Work
Despite MARE improving the accuracy and reliability of VLMs in Deepfake detection and reasoning, it also has several limitations that warrant further investigation. First, the DMA dataset relies on manual annotation for detailed forgery text descriptions, which limits its scalability. A promising direction for future work is to explore automated or semi-automated pipelines for constructing explainable Deepfake datasets. Besides, existing evaluation methodologies are becoming insufficient for assessing the logical consistency and explainability of explainable Deepfake detection. Establishing more effective evaluation methodologies, potentially based on information theory of distribution hypotheses, is a valuable research direction.